OpenStudy (anonymous):

You bought a TV with a sale price of $960. If the markdown price was 20% off the original price, what was the original price of the TV?

OpenStudy (anonymous):

plz help

OpenStudy (goldphenoix):

You can use a proportion to solve this.

OpenStudy (anonymous):

what kind?

OpenStudy (goldphenoix):

Actually, another method is better.

OpenStudy (anonymous):

(1.00 - 0.20)P = 960 (0.80)P = 960 P = 960/(0.80) = 1200

OpenStudy (goldphenoix):

^ That's what I was going to say.

OpenStudy (anonymous):

btw, where I derived (1.00 - 0.20) is 100% - 20%

OpenStudy (anonymous):

all good now @aprilsages ?

OpenStudy (goldphenoix):

You can go backward to solve this. Let say x is the cost of the original price. (100%). But the sale price is $960 after you cut 20 percent off the original price. So 100% - 20% = 80%. You're still paying 80 percent of the original price. So by using his equation, \[\large \large (80% \times OriginalPrice) = $960. \] Solve it, and that will give you the original price. Though he already gave you the answer $1200, so the original price is $1200.

OpenStudy (anonymous):

all good thanks @tcarroll010 @GoldPhenoix

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @aprilsages

OpenStudy (goldphenoix):

\[\large \large (80percent) \times (OriginalPrice) = 960\] Sorry, didn't know the percent sign didn't work.

OpenStudy (anonymous):

ty!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

uw!